Veronica has a garden in the shape of a rectangle that has an area of 78 yards. The length of the garden is 1 yard more than twice the width what are the length and width of veronicas garden
Answer by palanisamy(496)

Let the width = x
Then its length = 2x+1
Area is x(2x+1) = 78
2x^2+x-78 = 0
2x^2+13x-12x-78=0
x(2x+13)-6(2x+13) = 0
(2x+13)(x-6) = 0
x = -13/2 or x=6
x cannot be negative.
So x = 6
Therefpre the width = 6
Then its length = 2x+1 = 2*6+1 = 13 m